A március 17.-i valószínűségszámítás dolgozat eredménye

1.  2.  3.  4.  5.  6.

diákok:

Baksa Dóra  +  +  −  +  −  +
Deme Adrienn  +  +  −  +  −  +
File Csenge  +  +  −  +  −  +
Ignéczi Andrea  +  +  −  +  −  +
Kiss Tibor  +  +  −  +  +  +
Koreczki Bence  +  ±  −  +  −  +
Nagy Ágnes  +  +  −  +  +  +
Pinczés József  −  +  −  −  −  +
Rácz Zsuzsanna  +  +  −  +  −  +
Szabó Timea  +  +  −  +  ±  +
Szikszai Márton  ±  +  −  ±  ±  +
Takács Csaba  +  +  −  +  −  +
Ungi Gergely  ±  +  −  +  ±  +
Varha Syuzanna  +  +  −  ±  −  +